A narrow beam of white light enters the narrow edge of a rectangular pane of glass at an angle of 60° from the direction perpendicular to the edge. The pane is 0.50 m in width. Estimate how far apart the red and blue parts of the visible spectrum will be when the light leaves the glass, first if the pane is made from crown glass and then if it is made from dense flint glass. (.0034m, .0083m)
Answer #1

law food runsas Apply Shens nisin Goa 2 Sinoa Sinbo = 1.512 Sinos. on or = s.nl/ singo) ( 512 - 34.940 tan/349) ha h = 0.349

DAMU op - Sint (sinbol) - 29.60 tar (29.66) = 65 h=0.2847 m Ora sint sinboo) = 29.930 tan (28.95) = 0 - 0.2764 - n - 0 2847 -